a rectangle has a perimeter of 38 cm. three-fourths of its width is equal to one-fifth of its length. what is the measure of the length?
Answer by edjones(8007)

A) 2L+2w=38
B)
L/5 = 3w/4
4L=15w Multiply each side by 20 to eliminate fractions.
4L-15w=0
.
-4L-4w=-76 multiply A) by -2
4L-15w=0
-----------add
==>-19w=-76
w=4 cm
L=15*4/4
=15 cm
.
Check:
A) 2*15 + 2*4 = 38 true
B) 15/5=3*4/4 true
.
Ed